OpenStudy (anonymous):

given the formula, y=3x+4, enter the value of y when the value of x is -2

OpenStudy (anonymous):

Y= 3(2)+4 =6+4 =10

OpenStudy (mathstudent55):

x = -2, not x = 2 y = 3(-2) + 4 = -6 + 4 = -2

OpenStudy (anonymous):

thank you guys

OpenStudy (mathstudent55):

wlcm